OpenStudy (ciaraxcx):

Write a^2 * a^3 without exponents. How do I do this? Will medal and fan of course.

zepdrix (zepdrix):

\[\large\rm a^2=a\cdot a\]\[\large\rm a^3=a\cdot a \cdot a\]So then,\[\large\rm a^2\cdot a^3=(a\cdot a)\cdot(a\cdot a\cdot a)\]How many a's multiplying in total? :)

OpenStudy (ciaraxcx):

Ohh I thought the question was asking for a much more complicated answer, my bad! Thanks for the help.

zepdrix (zepdrix):

Ya the wording is confusing.. without exponents? I think they meant to say "as a single exponent" or something like that..